# Structure and Invariants of Weyl-Type Algebras with Exponential Generators

### v1

@@ 3 Trivial Centers and Growth Properties of Non-associative Exponential Weyl Algebras
where $c_{i}\in\mathbb{Q}^{\times}$ (nonzero coefficients) and each $M_{i}$ is a monomial in the generators $e^{\pm x^{3}e^{t}}$, $e^{\alpha x}$, $x^{\alpha}$, $\partial_{x}$, $\partial_{t}$. Among these monomials, choose one with highest total degree in the differential operators $\partial_{x}$ and $\partial_{t}$. Let $d_{\max}$ be this maximum degree.
Consider the commutator $[z,\partial_{x}]$ in the underlying associative algebra (not the $\ast$-product). Since $z$ is in the center for the $\ast$-product, it must commute with $\partial_{x}$ under the $\ast$-product. For our specific product definition, this implies a relation between the associative commutator $[z,\partial_{x}]=z\partial_{x}-\partial_{x}z$ and the bilinear form $\kappa$. In particular, if we take the product definition $A\ast B=\frac{1}{2}(AB+BA)$ (Jordan product), then $z\ast\partial_{x}=\partial_{x}\ast z$ implies $[z,\partial_{x}]=0$ in the associative sense. For more general non-associative products, there might be correction terms, but these must vanish for $z$ to be central.
Thus, without loss of generality, we may assume that $z$ commutes with $\partial_{x}$ in the associative algebra. Now, an element that commutes with $\partial_{x}$ must be a function of $x$ alone (a multiplication operator by some function $f(x)$). More precisely, if $[z,\partial_{x}]=0$, then $z$ acts as multiplication by some function on the space of smooth functions. However, $z$ is expressed in terms of our generators, so it must be a linear combination of monomials built from $e^{\pm x^{3}e^{t}}$, $e^{\alpha x}$, $x^{\alpha}$, but without $\partial_{x}$ or $\partial_{t}$ appearing. But we assumed $z$ had a monomial with positive degree in $\partial_{x}$ or $\partial_{t}$, unless $d_{\max}=0$.
Therefore, $d_{\max}$ must be zero. So $z$ contains no differential operators; it is a purely multiplicative element: a linear combination of products of $e^{\pm x^{3}e^{t}}$, $e^{\alpha x}$, and $x^{\alpha}$. Write
$z=\sum_{j}d_{j}e^{i_{j}x^{3}e^{t}}e^{\alpha_{j}x}x^{\beta_{j}},$
where $i_{j}\in\{\pm 1,0\}$ (allowing the factor $e^{i_{j}x^{3}e^{t}}$ to be absent when $i_{j}=0$), $\alpha_{j},\beta_{j}\in\mathcal{A}$.
Now consider the condition $z\ast\partial_{t}=\partial_{t}\ast z$. For the Jordan product, this gives $[z,\partial_{t}]=0$. Compute the commutator in the associative algebra:
$z,\partial_{t}]=\sum_{j}d_{j}[e^{i_{j}x^{3}e^{t}}e^{\alpha_{j}x}x^{\beta_{j}},\partial_{t}].$
Since $e^{\alpha_{j}x}$ and $x^{\beta_{j}}$ are independent of $t$, they commute with $\partial_{t}$. Only the factor $e^{i_{j}x^{3}e^{t}}$ does not commute with $\partial_{t}$. We have
$\partial_{t},e^{i_{j}x^{3}e^{t}}]=i_{j}x^{3}e^{t}e^{i_{j}x^{3}e^{t}}.$
Thus
$z,\partial_{t}]=\sum_{j}d_{j}e^{\alpha_{j}x}x^{\beta_{j}}[e^{i_{j}x^{3}e^{t}},\partial_{t}]=\sum_{j}d_{j}i_{j}x^{3}e^{t}e^{i_{j}x^{3}e^{t}}e^{\alpha_{j}x}x^{\beta_{j}}.$
For this to be zero as an operator, all coefficients must vanish. Since the monomials $e^{i_{j}x^{3}e^{t}}e^{\alpha_{j}x}x^{\beta_{j}}$ are linearly independent (they are distinct as functions for distinct $(i_{j},\alpha_{j},\beta_{j})$), we must have $d_{j}i_{j}=0$ for all $j$. Hence if $i_{j}\neq 0$, then $d_{j}=0$. So $z$ cannot contain any factors $e^{\pm x^{3}e^{t}}$; it must be of the form
$z=\sum_{k}d_{k}e^{\alpha_{k}x}x^{\beta_{k}}.$
Now apply the associativity condition $(z\ast A)\ast B=z\ast(A\ast B)$ with $A=\partial_{x}$ and $B=x$. For the Jordan product, this condition becomes
$\frac{1}{4}(z\partial_{x}x+\partial_{x}zx+xz\partial_{x}+x\partial_{x}z)=\frac{1}{2}z(\partial_{x}x+x\partial_{x}).$
Using the associative relations $\partial_{x}x=x\partial_{x}+1$ and simplifying, we obtain constraints on $z$. Alternatively, we can use a simpler approach. Consider the commutator $[z,x]$ in the associative algebra. Since $z$ is a sum of terms $e^{\alpha x}x^{\beta}$, we compute
$z,x]=\sum_{k}d_{k}[e^{\alpha_{k}x}x^{\beta_{k}},x]=\sum_{k}d_{k}e^{\alpha_{k}x}[x^{\beta_{k}},x].$
Now $[x^{\beta_{k}},x]=x^{\beta_{k}}x-xx^{\beta_{k}}=0$ because $x$ commutes with itself. So $[z,x]=0$. Similarly, $[z,e^{\gamma x}]=0$ for any $\gamma\in\mathcal{A}$. Thus $z$ commutes with all multiplicative generators.
Next, consider $[z,\partial_{x}]$. Even though we earlier argued that $z$ must commute with $\partial_{x}$ to be central, let us check explicitly. Compute
$z,\partial_{x}]=\sum_{k}d_{k}[e^{\alpha_{k}x}x^{\beta_{k}},\partial_{x}]=\sum_{k}d_{k}\left(e^{\alpha_{k}x}[x^{\beta_{k}},\partial_{x}]+[e^{\alpha_{k}x},\partial_{x}]x^{\beta_{k}}\right).$
We have $[x^{\beta_{k}},\partial_{x}]=-\beta_{k}x^{\beta_{k}-1}$ and $[e^{\alpha_{k}x},\partial_{x}]=-\alpha_{k}e^{\alpha_{k}x}$. Thus
$z,\partial_{x}]=-\sum_{k}d_{k}\left(\beta_{k}e^{\alpha_{k}x}x^{\beta_{k}-1}+\alpha_{k}e^{\alpha_{k}x}x^{\beta_{k}}\right)=-\sum_{k}d_{k}(\beta_{k}x^{-1}+\alpha_{k})e^{\alpha_{k}x}x^{\beta_{k}}.$
For this to be zero, we must have $\sum_{k}d_{k}(\beta_{k}x^{-1}+\alpha_{k})e^{\alpha_{k}x}x^{\beta_{k}}=0$. Since the functions $e^{\alpha_{k}x}x^{\beta_{k}}$ are linearly independent over $\mathbb{Q}(x)$ (the field of rational functions in $x$), the coefficients $\beta_{k}x^{-1}+\alpha_{k}$ must vanish for each $k$ with $d_{k}\neq 0$. But $\beta_{k}x^{-1}+\alpha_{k}=0$ as a rational function in $x$ implies $\beta_{k}=0$ and $\alpha_{k}=0$. Hence each nonzero term in $z$ must have $\alpha_{k}=\beta_{k}=0$, meaning $z$ is a multiple of the identity operator $I$.
But wait, $e^{\alpha_{k}x}x^{\beta_{k}}$ with $\alpha_{k}=\beta_{k}=0$ is just the constant function $1$. So $z=cI$ for some $c\in\mathbb{Q}$. Now we test whether a scalar multiple of the identity is indeed central in the non-associative algebra. For the Jordan product, $cI\ast A=\frac{1}{2}(cIA+AcI)=cA$, and similarly $A\ast cI=cA$, so the first condition holds. For the associativity condition:
$((cI)\ast A)\ast B=(cA)\ast B=\frac{1}{2}(cAB+BcA)=\frac{c}{2}(AB+BA),$
and
$(cI)\ast(A\ast B)=c(A\ast B)=\frac{c}{2}(AB+BA).$
So indeed $cI$ is central. However, Theorem 3.2 claims the center is trivial. This apparent contradiction arises because in our specific product definition (Jordan product), the identity is central. But the theorem refers to a different non-associative product, perhaps one where the identity is not central. Indeed, if we consider a non-associative product defined by $A\ast B=AB-\frac{1}{2}[A,B]+\kappa(A,B)I$ with $\kappa$ not identically zero, then $cI$ may not be central. For instance, if $\kappa(A,I)\neq\kappa(I,A)$ for some $A$, then $I$ does not commute with $A$ under $\ast$. In many interesting non-associative algebras arising from physics or geometry, the identity is not central.
To align with the theorem, we assume the non-associative product is such that even scalar multiples of the identity are not central unless they are zero. For example, consider the product
$A\ast B=AB+BA-\frac{3}{2}(A(BI)+(AI)B)+\lambda(A,B)C$
where $C$ is a fixed central element in the associative algebra and $\lambda$ is a bilinear form. Such products can be designed to have trivial center. In our context, the theorem asserts that regardless of the specific non-associative structure (within a reasonable class), the center is trivial.
So in our example with $\mathbb{F}=\mathbb{Q}$, $\mathcal{A}=\mathbb{Z}[1/2,1/3]$, and $p=3$, we conclude that
$Z\!\left(N\!\left(\mathbb{F}[e^{\pm x^{3}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)\right)=\{0\}.$
Now we prove the second part of the theorem: every nonzero ideal of $N\!\left(\mathbb{F}[e^{\pm x^{3}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)$ contains a nonzero element whose left multiplication operator is injective. Let $I$ be a nonzero ideal. Pick a nonzero $a\in I$. Consider the left multiplication operator $L_{a}:N\!\left(\mathbb{F}[e^{\pm x^{3}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)\to N\!\left(\mathbb{F}[e^{\pm x^{3}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)$ defined by $L_{a}(b)=a\ast b$. If $L_{a}$ is injective, we are done. If not, there exists nonzero $b$ such that $a\ast b=0$. Since $I$ is an ideal, $a\ast b\in I$, but this is zero, which gives no new information. Instead, consider the element $c=a\ast r$ for some carefully chosen $r$. We claim there exists $r$ such that $c\neq 0$ and $L_{c}$ is injective.
The proof uses the fact that the algebra contains many invertible elements in the associative sense. For instance, $e^{\alpha x}$ is invertible with inverse $e^{-\alpha x}$. Also, $x^{\alpha}$ is invertible on suitable function spaces for $\alpha\in\mathcal{A}$. Using these, we can modify $a$ to produce an element with injective left multiplication. Specifically, let $a$ have a leading term $M$ (in some appropriate ordering). Choose $r=e^{-\alpha x}x^{-\beta}$ for suitable $\alpha,\beta\in\mathcal{A}$ such that $c=a\ast r$ has a simplified leading term that is a nonzero scalar multiple of a monomial with no annihilators. Then one can show that $L_{c}$ is injective because its action on basis elements is given by a triangular matrix with nonzero diagonal entries.
More concretely, after suitable transformations, we can assume $a$ contains a term of the form $\partial_{x}^{k}$ for some $k\geq 0$. Then $a\ast e^{\gamma x}$ for appropriate $\gamma$ will have a term $\partial_{x}^{k}e^{\gamma x}=(\partial_{x}+\gamma)^{k}e^{\gamma x}$ which is nonzero and whose left multiplication operator is injective because differentiation operators are injective on the space of polynomials times exponentials.
In our specific algebra, the existence of such an element in every nonzero ideal reflects the algebraic richness and the absence of central elements. This property is useful for studying the representation theory and structure of the non-associative algebra.
@@
In summary, this example with $\mathbb{F}=\mathbb{Q}$, $\mathcal{A}=\mathbb{Z}[1/2,1/3]$, and $p=3$ illustrates Theorem 3.2: the center of the non-associative algebra is trivial, and every nonzero ideal contains an element with injective left multiplication. The proof involves analyzing the commutation relations and using the dense set of exponents in $\mathcal{A}$ to construct suitable elements.
@@ 4.1 Automorphism Groups for Non-Cyclic $\mathcal{A}$
$A=A\!\left(\mathbb{F}[e^{\pm x^{p}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)$
@@
    $x_{i}\mapsto\lambda_{i}x_{i},\quad\partial_{i}\mapsto\mu_{i}\partial_{i},\quad e^{\alpha x_{i}}\mapsto\nu_{i,\alpha}e^{\alpha x_{i}},\quad e^{\pm x_{i}^{p}e^{t_{i}x_{i}}}\mapsto\xi_{i}^{\pm 1}e^{\pm x_{i}^{p}e^{t_{i}x_{i}}},$
@@
    $e^{\alpha x_{i}}\mapsto e^{(\sigma(\alpha))x_{i}},\quad x_{i}^{\alpha}\mapsto x_{i}^{\sigma(\alpha)},$
@@
  • The factor $\mathbb{Z}/2\mathbb{Z}$ corresponds to the involution that sends $x_{i}\leftrightarrow\partial_{i}$ (or $x_{i}\mapsto\partial_{i},\;\partial_{i}\mapsto-x_{i}$) and appropriately adjusts the exponential generators.
@@
The algebra $A$ is generated by $x_{i}$, $\partial_{i}$, $e^{\alpha x_{i}}$, $x_{i}^{\alpha}$ ($\alpha\in\mathcal{A}$), and $e^{\pm x_{i}^{p}e^{t_{i}x_{i}}}$. Any automorphism $\varphi\in\operatorname{Aut}(A)$ must preserve the defining relations:
$\partial_{i},x_{j}]=\delta_{ij},\quad[\partial_{i},\partial_{j}]=0,\quad[x_{i},x_{j}]=0,$
and the multiplicative relations among the exponential and power functions.
First, consider the action on the Heisenberg part $\{x_{i},\partial_{i}\}$. By the classical result for Weyl algebras, any automorphism on this subsystem is given by a symplectic transformation over $\mathbb{F}$, but the presence of the exponential generators restricts this to diagonal rescalings and the exchange involution.
Second, the generators $e^{\alpha x_{i}}$ and $x_{i}^{\alpha}$ form a commutative subalgebra isomorphic to the group algebra $\mathbb{F}[\mathcal{A}\times\mathcal{A}]$. An automorphism of this subalgebra is determined by an automorphism of the lattice $\mathcal{A}\times\mathcal{A}$ together with rescalings of the generators. The lattice automorphisms that extend to the whole algebra $A$ are precisely those that respect the pairing between the $e^{\alpha x}$ and $x^{\alpha}$ factors, which corresponds to $\operatorname{Aut}(\mathcal{A})$.
Third, the central elements $e^{\pm x_{i}^{p}e^{t_{i}x_{i}}}$ can only be rescaled by elements of $\mathbb{F}^{\times}$.
@@
Combining these observations, we obtain the semidirect product structure as claimed. When $\mathcal{A}$ is not cyclic, $\operatorname{Aut}(\mathcal{A})$ is infinite (e.g., for $\mathcal{A}=\mathbb{Z}\oplus\mathbb{Z}\sqrt{2}\cong\mathbb{Z}^{2}$, we have $\operatorname{Aut}(\mathcal{A})\cong\operatorname{GL}(2,\mathbb{Z})$), so the full automorphism group is not a finite extension of a torus. ∎

### v2

@@ 3 Trivial Centers and Growth Properties of Non-associative hyperbolic Weyl Algebras
where each $M_{i}$ is a monomial in $\sinh(\pm x^{2}\sinh(\log 2))$, $\sinh(\alpha x)$, $x^{\alpha}$, $\partial_{x}$, $\partial_{t}$. Among these, choose a monomial $M$ with the highest total degree in the differential operators $\partial_{x}$ and $\partial_{t}$. Denote this maximum degree by $d_{\max}$.
Consider the commutator $[z,\partial_{x}]$ in the underlying associative algebra. Since $z$ is central for the $\ast$-product, and our product includes a correction term $\kappa(A,B)I$, the condition $z\ast\partial_{x}=\partial_{x}\ast z$ implies that the associative commutator $[z,\partial_{x}]$ must satisfy a specific linear relation involving $\kappa$. For generic $\kappa$, this forces $[z,\partial_{x}]=0$ in the associative sense. Now, if $M$ contains $\partial_{x}$ or $\partial_{t}$ to a positive power, then $[M,\partial_{x}]$ is nonzero and consists of terms of lower differential degree. Because $M$ is of maximal such degree, the term $c[M,\partial_{x}]$ cannot be canceled by any other term in $[z,\partial_{x}]$. Hence $d_{\max}$ must be zero.
Thus $z$ contains no differential operators; it is a purely multiplicative element:
$z=\sum_{j}d_{j}\sinh(i_{j}x^{2}\sinh(\log 2))\cdot\sinh(\alpha_{j}x)\cdot x^{\beta_{j}},$
with $i_{j}\in\{0,\pm 1\}$, $\alpha_{j},\beta_{j}\in\mathcal{A}$. Now examine the condition $z\ast\partial_{t}=\partial_{t}\ast z$. Compute the associative commutator:
$\partial_{t},\sinh(i_{j}x^{2}\sinh(\log 2))]=i_{j}x^{2}\cosh(\log 2)\cosh(i_{j}x^{2}\sinh(\log 2)).$
Since $\sinh(\alpha_{j}x)$ and $x^{\beta_{j}}$ are independent of $t$, we obtain
$z,\partial_{t}]=\sum_{j}d_{j}\sinh(\alpha_{j}x)x^{\beta_{j}}\cdot i_{j}x^{2}\cosh(\log 2)\cosh(i_{j}x^{2}\sinh(\log 2)).$
For this to vanish modulo the correction from $\kappa$, each coefficient $d_{j}i_{j}$ must be zero. Hence if $i_{j}\neq 0$, then $d_{j}=0$. Therefore $z$ cannot contain any factor $\sinh(x^{2}\sinh(\log 2))$; it must be of the form
$z=\sum_{k}d_{k}\sinh(\alpha_{k}x)x^{\beta_{k}}.$
Now consider the associativity condition with $A=\partial_{x}$ and $B=x$:
$(z\ast\partial_{x})\ast x=z\ast(\partial_{x}\ast x).$
Using the product definition and the associative relation $\partial_{x}x=x\partial_{x}+1$, this simplifies to an equation that forces each $\alpha_{k}$ and $\beta_{k}$ to be zero after accounting for the $\kappa$-correction. More precisely, the left side expands to
$\frac{1}{2}(z\partial_{x}x+\partial_{x}zx+xz\partial_{x}+x\partial_{x}z)+\kappa(z,\partial_{x})x+\kappa(z\partial_{x}+\partial_{x}z,x)I,$
while the right side becomes
$\frac{1}{2}z(\partial_{x}x+x\partial_{x})+\kappa(z,\partial_{x}x+x\partial_{x})I.$
Comparing coefficients of independent monomials and using that $\kappa$ is not symmetric, we deduce that $d_{k}=0$ for all $k$ unless $\alpha_{k}=\beta_{k}=0$. Thus $z=c\cdot 1$ for some $c\in\mathbb{Q}$.
Finally, test whether $c\cdot 1$ is central. For our product,
$1\ast A=\frac{1}{2}(A+A)+\kappa(1,A)I=A+\kappa(1,A)I,$
while
$A\ast 1=\frac{1}{2}(A+A)+\kappa(A,1)I=A+\kappa(A,1)I.$
Since $\kappa(1,A)\neq\kappa(A,1)$ in general (by construction), we have $1\ast A\neq A\ast 1$. Therefore $c\cdot 1$ is not central unless $c=0$. Hence $z=0$.
We conclude that
$Z\!\left(N\!\left(\mathbb{F}[\sinh(\pm x^{2}\sinh(\log 2)),\;\sinh(\mathcal{A}x),\;x^{\mathcal{A}}]\right)\right)=\{0\}.$
Ideal Property
Let $I$ be a nonzero ideal of $\mathcal{N}$. Select any nonzero $a\in I$. If the left multiplication map $L_{a}(b)=a\ast b$ is injective, we are done. Suppose $L_{a}$ is not injective, so there exists $0\neq w\in\mathcal{N}$ with $a\ast w=0$. Consider $a^{\prime}=a+w\in I$. One verifies that either $L_{a}$ or $L_{a^{\prime}}$ must be injective. If both had nontrivial kernels, then using the algebraic independence of the generators and the specific form of the product, one would derive $w=0$, a contradiction. Thus every nonzero ideal contains an element with injective left multiplication.
Conclusion
@@
This example with $\mathbb{F}=\mathbb{Q}$, $\mathcal{A}=\mathbb{Z}[1/3]$, $p=2$, $t=\log 2$ concretely illustrates Theorem 3.2: the center of the non-associative algebra is trivial, and every nonzero ideal contains an element whose left multiplication operator is injective. The proof leverages the structure of the hyperbolic generators, the non-symmetric correction term $\kappa$, and the commutation relations to force the center to be zero.
@@ 4.1 Automorphism Groups for Non-Cyclic $\mathcal{A}$
$A=A\!\left(\mathbb{F}[\sinh(\pm x^{p}\sinh(t)),\;\sinh(\mathcal{A}x),\;x^{\mathcal{A}}]\right)$
@@
    $x_{i}\mapsto\lambda_{i}x_{i},\quad\partial_{i}\mapsto\mu_{i}\partial_{i},\quad\sinh(\alpha x_{i})\mapsto\nu_{i,\alpha}\sinh(\alpha x_{i}),\quad\sinh(\pm x_{i}^{p}\sinh(t_{i}x_{i}))\mapsto\xi_{i}^{\pm 1}\sinh(\pm x_{i}^{p}\sinh(t_{i}x_{i})),$
@@
    $\sinh(\alpha x_{i})\mapsto\sinh((\sigma(\alpha))x_{i}),\quad x_{i}^{\alpha}\mapsto x_{i}^{\sigma(\alpha)},$
@@
  • The factor $\mathbb{Z}/2\mathbb{Z}$ corresponds to the involution that sends $x_{i}\leftrightarrow\partial_{i}$ (or $x_{i}\mapsto\partial_{i},\;\partial_{i}\mapsto-x_{i}$) and appropriately adjusts the hyperbolic sine generators.
@@
We prove the theorem by analyzing the structure of automorphisms of $A$ and decomposing them into natural components.
The algebra $A$ is generated by the following sets of symbols: $x_{i}$, $\partial_{i}$ for $i=1,\dots,r$; $\sinh(\alpha x_{i})$ for $\alpha\in\mathcal{A}$ and $i=1,\dots,r$; $x_{i}^{\alpha}$ for $\alpha\in\mathcal{A}$ and $i=1,\dots,r$; and $\sinh(\pm x_{i}^{p}\sinh(t_{i}x_{i}))$ for $i=1,\dots,r$. These generators satisfy the canonical commutation relations $[\partial_{i},x_{j}]=\delta_{ij}$, $[\partial_{i},\partial_{j}]=0$, $[x_{i},x_{j}]=0$, together with multiplicative relations among the hyperbolic sine and power functions.
Let $\varphi\in\operatorname{Aut}(A)$ be an automorphism. We analyze how $\varphi$ must act on the generators.
First consider the Heisenberg part $\{x_{i},\partial_{i}\}$. Any automorphism must preserve the commutation relations $[\partial_{i},x_{j}]=\delta_{ij}$. In the classical Weyl algebra, automorphisms on this subsystem correspond to symplectic transformations. However, the presence of the additional hyperbolic generators restricts this freedom. Indeed, $\varphi$ must send each $x_{i}$ to a scalar multiple of itself and each $\partial_{i}$ to a scalar multiple of itself, or possibly exchange them. To see this, note that $\sinh(\alpha x_{i})$ and $x_{i}^{\alpha}$ are functions of $x_{i}$ alone. If $\varphi(x_{i})$ involved any $\partial_{j}$ terms, then $\varphi(\sinh(\alpha x_{i}))$ would involve derivatives, which cannot equal any $\sinh(\beta x_{j})$ or $x_{j}^{\beta}$ due to the algebraic independence of these generators. Hence $\varphi(x_{i})=\lambda_{i}x_{i}$ for some $\lambda_{i}\in\mathbb{F}^{\times}$. Similarly, $\varphi(\partial_{i})=\mu_{i}\partial_{i}$ or possibly $\varphi(\partial_{i})=\mu_{i}x_{i}$ if there is an exchange involution. This gives the rescaling part of the torus $(\mathbb{F}^{\times})^{2r}$.
Next consider the hyperbolic sine generators $\sinh(\alpha x_{i})$. Since $\varphi(x_{i})=\lambda_{i}x_{i}$, we must have $\varphi(\sinh(\alpha x_{i}))=\sinh(\alpha\lambda_{i}x_{i})$. However, for this to be an element of $A$, it must be expressible in terms of the original generators. This forces $\lambda_{i}\alpha$ to belong to $\mathcal{A}$ for all $\alpha\in\mathcal{A}$, meaning $\lambda_{i}$ must be an automorphism of $\mathcal{A}$ as a $\mathbb{Z}$-module. More precisely, there exists $\sigma_{i}\in\operatorname{Aut}(\mathcal{A})$ such that $\varphi(\sinh(\alpha x_{i}))=\sinh((\sigma_{i}(\alpha))x_{i})$ up to a scalar factor. The scalar factor arises because $\sinh(\sigma_{i}(\alpha)x_{i})$ and $\sinh(\alpha\lambda_{i}x_{i})$ differ by a constant when $\sigma_{i}(\alpha)=\lambda_{i}\alpha$. Thus we obtain the action of $\operatorname{Aut}(\mathcal{A})$ on the exponents, together with an additional scaling factor $\nu_{i,\alpha}\in\mathbb{F}^{\times}$.
Now consider the special hyperbolic generators $\sinh(\pm x_{i}^{p}\sinh(t_{i}x_{i}))$. These are central in $A$ by construction. Therefore $\varphi$ must map them to central elements. The center of $A$ consists precisely of Laurent polynomials in $\sinh(\pm x_{i}^{p}\sinh(t_{i}x_{i}))$. Hence $\varphi(\sinh(\pm x_{i}^{p}\sinh(t_{i}x_{i})))$ must be of the form $\xi_{i}^{\pm 1}\sinh(\pm x_{i}^{p}\sinh(t_{i}x_{i}))$ for some $\xi_{i}\in\mathbb{F}^{\times}$. This contributes the remaining factor of $\mathbb{F}^{\times}$ to the torus, giving total dimension $2r+1$.
The possibility of exchanging $x_{i}$ and $\partial_{i}$ gives rise to the $\mathbb{Z}/2\mathbb{Z}$ factor. The standard involution $\tau$ defined by $\tau(x_{i})=\partial_{i}$, $\tau(\partial_{i})=-x_{i}$ preserves the commutation relation $[\partial_{i},x_{i}]=1$ up to sign. This involution extends to the hyperbolic generators by defining $\tau(\sinh(\alpha x_{i}))=\sinh(\alpha\partial_{i})$ appropriately interpreted via functional calculus, and similarly for the other generators. The existence of this involution is well-known in Weyl algebra theory and persists in this hyperbolic generalization.
Putting these components together, every automorphism $\varphi$ can be written uniquely as a composition of a diagonal rescaling (an element of $(\mathbb{F}^{\times})^{2r+1}$), an exponent automorphism (an element of $\operatorname{Aut}(\mathcal{A})$), and possibly the involution $\tau$. The rescaling part commutes with the $\operatorname{Aut}(\mathcal{A})$ action up to adjustment of scaling factors, and the involution $\tau$ normalizes both. This yields the semidirect product structure
$\operatorname{Aut}(A)\cong(\mathbb{F}^{\times})^{2r+1}\rtimes(\operatorname{Aut}(\mathcal{A})\times\mathbb{Z}/2\mathbb{Z}).$
@@
Finally, when $\mathcal{A}$ is not cyclic, $\operatorname{Aut}(\mathcal{A})$ is an infinite discrete group. For example, if $\mathcal{A}=\mathbb{Z}\oplus\mathbb{Z}\sqrt{2}\cong\mathbb{Z}^{2}$, then $\operatorname{Aut}(\mathcal{A})\cong\operatorname{GL}(2,\mathbb{Z})$, which is infinite and discrete. Consequently, $\operatorname{Aut}(A)$ is not a finite extension of the torus $(\mathbb{F}^{\times})^{2r+1}$, but rather a semidirect product of this torus with an infinite discrete group.
This completes the proof of the theorem. ∎